# Enhancing Semantic Segmentation with Continual Self-Supervised Pre-training

### v1

@@ 4.2 Regional level consistency
  • Attention–aware region sampling: the attention map of a block of a self-supervised ViT encoder often contain several insights on an image. In fact, over the different heads of the last block, the attention is directed towards different regions, as presented in Figure˜2. Consequently, to enforce more semantically-rich regions, we use the attention from the different heads of the encoder to generate the starting patch for candidate regions. In this case, a starting patch is the one getting the most attention on a specific head. From the starting patch, we define the region using a similar process as in random sampling. This is feasible primarily in the context of continual pre-training, as it leverages publicly available pre-trained models that already exhibit a useful signal for attention-aware region sampling—something not achievable when training models from scratch.
@@
$\tilde{z}_{r}=\operatorname{CA}(z_{r},Z,Z)=(W_{v}Z)\textnormal{softmax}(\tau(W_{k}Z)^{T}(W_{q}z_{r})),$ (5)
@@
$\tilde{z}^{\prime}_{r}=\operatorname{CA}(z^{\prime}_{r},Z_{R},R)=(W_{v}Z_{R})\textnormal{softmax}(\tau(W_{k}Z_{R})^{T}(W_{q}z^{\prime}_{r})),$ (6)
@@ 4.3 Local consistency
Firstly, we aim to alleviate the problem of limited data via separate patch-level augmentations for a fraction of the total image patches. Specifically, we consider strong blurring on 30% of random patches on the base encoder views. This process creates incomplete information on the augmented patches. However, unlike the standard practice in Masked Image Modeling (MIM) (Zhou et al., 2022; Zhang et al., 2023a; Oquab et al., 2024), we do not feed empty tokens for the blurred patches. Thus, we expect the following: i) the attention between the local patches in the ViT model are used to “complete” the local patch feature and ii) additional scale robustness is imposed on the image patch feature via inter-patch consistency. The main reason for not using MIM in this case, especially for out-of-domain datasets, is that it is known to be data-intensive, requiring a large number of samples and training time to learn meaningful representations (He et al., 2022). This challenge becomes even more pronounced with new domains that exhibit high variability, unlike standard datasets like ImageNet (Deng et al., 2009), and with few samples as typically encountered in continual pre-training. We also investigate other patch-level augmentations such as rotation and noising, and blurring turns out to be the best-performing approach.
@@
5 Experiment Setup
We use ViT-S/16 (Dosovitskiy et al., 2020) for our experiments due to its cost-efficiency balance between training cost and performance, as in Su & Ji (2024). We apply continual pre-training for 100 epochs on the target dataset, after observing lower performance when training longer. We also used a register token in our ViT encoder as in Darcet et al. (2024); Su & Ji (2024). We use a batch size of 512, and a shared projection head across the different pre-training levels as done in Su & Ji (2024); Zhang et al. (2023b); Zhou et al. (2022) with the output dimension of $K=8192$. The learning rate is linearly increased for the first epoch to its base value calculated using the linear scaling rule in Chen et al. (2020b), which is $lr=1.5\cdot 10^{-4}$. After warmup, the learning rate is decreased using a cosine scheduler (Loshchilov & Hutter, 2022). The weight decay is set to 0.1 and we use AdamW optimizer (Loshchilov & Hutter, 2017). We follow the data augmentations of BYOL (Grill et al., 2020) which was also used in Caron et al. (2021) (i.e. color jittering, gaussian blur, solarization) on random resized crops. Specifically, given an input image we generate 2 global views of resolution $224\times 224$ and 10 local views of resolution $96\times 96$ as in DINO (Caron et al., 2021).
We compare GLARE with SOTA SSL methods described in Su et al. (2023); Su & Ji (2024) on ViT-S by analyzing the performances of these pre-training methods against GLARE when they are used from scratch or in a continual pre-training setup. In this work, we initialize our model using the pre-trained weights from UDI (Su & Ji, 2024), which currently represents the state-of-the-art in leveraging self-supervised learning for downstream segmentation tasks. We then train only the parameters of the adapter layer as shown in Section˜3.2.
@@ 6.1 Semantic segmentation performance on different benchmarks
In Table˜1, we report the performance on semantic segmentation of models obtained by continual pre-training with different methods starting from UDI weights (Su & Ji, 2024). We also include the performance when the model is randomly initialized without any pre-training knowledge. We report the average mIoU scores and standard deviation after finetuning $3\times$ the pre-trained models on semantic segmentation using FPN (Kirillov et al., 2019). We consider 3 classes of datasets, namely general domain: ADE20k (Zhou et al., 2017), Pascal Context (Mottaghi et al., 2014), driving: Cityscapes (Cordts et al., 2016) and aerial: LoveDA (Wang et al., 2021a), which contain respectively $20k$, $4998$, $2975$ and $2522$ images in their training set, making them suitable for continual pre-training in limited data scenarios (<=20k<=20k images). For each experiment, we first initialize the weights of the backbone with those of the starting model (UDI), then we do continual SSL pre-training with an adapter (UniAdapter) on the target dataset by training only the adapter parameters. The obtained model is then used for finetuning for segmentation. We observe that GLARE reports consistent improvement for continual pre-training from UDI weights over all the datasets. On ADE20k, we obtain an improvement of +0.4+0.4 over UDI and on LoveDA, which is an out-of-domain dataset with respect to the original pre-training dataset of UDI (ImageNet (Deng et al., 2009)), we achieve an improvement of +0.6+0.6. This shows that GLARE is able to take advantage of existing encoded features and new data distribution to improve semantic understanding, which transfer in semantic segmentation.
Table 1: Comparison of SSL pre-trained models and continual pre-trained models starting from UDI (Su & Ji, 2024) on 4 semantic segmentation benchmarks. We report mIoU on the validation sets. We use the FPN (Kirillov et al., 2019) framework with 20k iterations and 2k for LoveDA. GLARE continual pre-training from UDI consistently shows improvements over the other pre-training strategies.
Method Backbone ADE20k P.Cont Cityscapes LoveDA
@@ 6.2 Comparison with other continual pre-training methods
Table 2: Comparison of different continual pre-training framework. Experiments performed starting from UDI (Su & Ji, 2024) pre-trained on ImageNet (Deng et al., 2009). We show the segmentation performance of continual pre-trained models on LoveDA with a finetuning for 2k iterations.
@@ 6.3 Influence of the SSL backbone
An important question is the role of the SSL backbone used for the continual pre-training of GLARE. Specifically, we want to examine whether the performance gains observed when starting from UDI also hold when initializing from other SSL backbones, in particular FLSL (Su et al., 2023) and DINO (Caron et al., 2021), all pre-trained on ImageNet-1k (Deng et al., 2009). To investigate this, Table˜3 reports the results of continual pre-training with GLARE starting from these different backbones, using both HPT (Reed et al., 2022) and UniAdapter (Lu et al., 2023) frameworks. We use LoveDA (Wang et al., 2021a) as the target domain and evaluate performance after fine-tuning for semantic segmentation. We observe that GLARE consistently improves upon the original models. Moreover, starting from a stronger backbone (e.g., UDI) leads to larger gains. We further hypothesize that the original pre-training strategy influences how effectively GLARE can leverage a given backbone.
@@ 6.4 Do the continual pre-trained models forget?
Pre-training Ptr. Data ADE20k P.Cont Citys.
UDI (Su & Ji, 2024) ImageNet (Deng et al., 2009) 41.1 49.2 74.7
@@ 6.5 Results in Classification
While our continual pre-training framework is primarily evaluated on semantic segmentation, it is not limited to this task. To demonstrate its broader applicability, we also apply GLARE continual pre-training to classification. Specifically, we experiment with two out-of-domain datasets: Derm7pt (Kawahara et al., 2019), a dermoscopic benchmark for skin lesion analysis containing $\sim$800 samples, and COVIDx (Wu et al., 2023), a large-scale x-ray benchmark for COVID-19 detection, where we subsample $20\%$ of the data for continual pre-training and fine-tuning to maintain a low-data setting. Models are initialized either randomly, from a DINO (Caron et al., 2021) pre-trained backbone, or from GLARE continual pre-training starting from DINO, and finally finetuning of a classification head. We report in Table˜5 the top-1 accuracy of models trained from scratch, initialized from DINO, or from DINO $\rightarrow$ GLARE. We observe that DINO $\rightarrow$ GLARE consistently outperforms the other initializations, demonstrating the effectiveness of GLARE continual pre-training even for tasks beyond semantic segmentation.
@@ 7.1 Different levels of understanding in GLARE
Table 6: Impact of the different pre-training levels in GLARE. We report the mIoU of finetuned continual pre-trained models when trained with different levels of GLARE on $20\%$ of ADE20k
@@ Appendix A Details about Datasets
In this paper, we perform continual learning and report results on the 4 semantic segmentation datasets: ADE20k (Zhou et al., 2017), Pascal Context (P.Cont) (Mottaghi et al., 2014), Cityscapes (Citys.) (Cordts et al., 2016) and LoveDA (Wang et al., 2021a). We use the training set of these datasets to apply our continual learning setup with different pre-training methods such as UDI (Su & Ji, 2024), FLSL (Su et al., 2023), GLARE. We then finetune the resulting models for segmentation using FPN (Kirillov et al., 2019) and report the performances of the models on the validation sets. In the following we provide more details on these datasets.
@@
LoveDA (Wang et al., 2021a). This focuses on different geographical environments between urban and rural. It contains high spatial resolution (0.3m) remote sensing images containing objects at different scales, complex backgrounds and inconsistent class distributions. Its training set is composed of 2,522 images, its validation set is composed of 1,669 images and test set of 1,796 images.
@@ C.1 Ablation of region sampling
Table 9: Ablation of the region sampling strategy. Experiments performed on UDI-GLARE with a finetuning of 2k iterations on LoveDA. $M$ represents the number of sampled regions considered.
@@ C.2 Effect of blurring strategy
random init. 19.3 37.4 34.4
block-wise 51.5 70.2 64.1
random 51.5 70.2 64.3
@@ Appendix D Visualization of Attention Maps
In this section we visualize some attention maps from the last block of the ViT encoder, using the [CLS] token as the query token. Figure˜4 shows the attention from DINO (Caron et al., 2021), the original pre-trained weights from UDI, and a GLARE continual pre-trained model starting from UDI (Su & Ji, 2024) weights, finetuned on ADE20k. We observe similarities in how the attention is distributed across the images, focusing on various details such as foreground objects, object parts, and the background. GLARE continual pre-trained model demonstrates reduced noise relative to UDI, with its attention more precisely directed toward specific objects or regions. When using GLARE for continual pre-training, the model leverages what has been learned before and learns supplementary semantics specific to the dataset.

### v2

@@ 4.2 Regional level consistency
  • Attention–aware region sampling: the attention map of a block of a self-supervised ViT encoder often contain several insights on an image. In fact, over the different heads of the last block, the attention is directed towards different regions, as presented in Figure˜2. Consequently, to enforce more semantically-rich regions, we use the attention from the different heads of the encoder to generate the starting patch for candidate regions. In this case, a starting patch is the one getting the most attention on a specific head. From the starting patch, we define the region using a similar process as in random sampling.
    This is feasible primarily in the context of continual pre-training, as it leverages publicly available pre-trained models that already exhibit a useful signal for attention-aware region sampling—something not achievable when training models from scratch.
@@
$\tilde{z}_{r}=\operatorname{CA}(z_{r},Z,Z)=(W_{v}Z)\operatorname{softmax}(\tau(W_{k}Z)^{T}(W_{q}z_{r})),$ (5)
@@
$\tilde{z}^{\prime}_{r}=\operatorname{CA}(z^{\prime}_{r},Z_{R},R)=(W_{v}Z_{R})\operatorname{softmax}(\tau(W_{k}Z_{R})^{T}(W_{q}z^{\prime}_{r})),$ (6)
@@ 4.3 Local consistency
Firstly, we aim to alleviate the problem of limited data via separate patch-level augmentations for a fraction of the total image patches. Specifically, we consider strong blurring on 30% of random patches on the base encoder views. This process creates incomplete information on the augmented patches. However, unlike the standard practice in Masked Image Modeling (MIM) (Zhou et al., 2022; Zhang et al., 2023a; Oquab et al., 2024), we do not feed empty tokens for the blurred patches. Thus, we expect the following: i) the attention between the local patches in the ViT model are used to “complete” the local patch feature and ii) additional scale robustness is imposed on the image patch feature via inter-patch consistency. The main reason for not using MIM in this case, especially for out-of-domain datasets, is that it is known to be data-intensive, requiring a large number of samples and training time to learn meaningful representations (He et al., 2022). This challenge becomes even more pronounced with new domains that exhibit high variability, unlike standard datasets like ImageNet (Deng et al., 2009), and with few samples as typically encountered in continual pre-training.
We also investigate other patch-level augmentations such as rotation and noising, and blurring turns out to be the best-performing approach.
@@
5 Experimental Setup
We use ViT-S/16 (Dosovitskiy et al., 2021) for our experiments due to its cost-efficiency balance between training cost and performance, as in Su and Ji (2024). We apply continual pre-training for 100 epochs on the target dataset, after observing lower performance when training longer. We also used a register token in our ViT encoder as in Darcet et al. (2024); Su and Ji (2024). We use a batch size of 512, and a shared projection head across the different pre-training levels as done in Su and Ji (2024); Zhang et al. (2023b); Zhou et al. (2022) with the output dimension of $K=8192$. The learning rate is linearly increased for the first epoch to its base value calculated using the linear scaling rule in Chen et al. (2020b), which is $lr=1.5\cdot 10^{-4}$. After warmup, the learning rate is decreased using a cosine scheduler (Loshchilov and Hutter, 2017). The weight decay is set to 0.1 and we use AdamW optimizer (Loshchilov and Hutter, 2019). We follow the data augmentations of BYOL (Grill et al., 2020) which was also used in Caron et al. (2021) (i.e. color jittering, gaussian blur, solarization) on random resized crops. Specifically, given an input image we generate 2 global views of resolution $224\times 224$ and 10 local views of resolution $96\times 96$ as in DINO (Caron et al., 2021).
We compare GLARE with SOTA SSL methods described in Su et al. (2023); Su and Ji (2024) on ViT-S by analyzing the performances of these pre-training methods against GLARE when they are used from scratch or in a continual pre-training setup. In this work, we initialize our model using the pre-trained weights from UDI (Su and Ji, 2024), which currently represents the state-of-the-art in leveraging self-supervised learning for downstream segmentation tasks. We then train only the parameters of the adapter layer as shown in Section˜3.2.
@@ 6.1 Semantic segmentation performance on different benchmarks
In Table˜1, we report the performance on semantic segmentation of models obtained by continual pre-training with different methods starting from UDI weights (Su and Ji, 2024). We also include the performance when the model is randomly initialized without any pre-training knowledge. We report the average mIoU scores and standard deviation after finetuning $3\times$ the pre-trained models on semantic segmentation using FPN (Kirillov et al., 2019). We consider 3 classes of datasets, namely general domain: ADE20k (Zhou et al., 2017), Pascal Context (Mottaghi et al., 2014), driving: Cityscapes (Cordts et al., 2016) and aerial: LoveDA (Wang et al., 2021a), which contain respectively $20k$, $4998$, $2975$ and $2522$ images in their training set, making them suitable for continual pre-training in limited data scenarios (<=20k<=20k images). For each experiment, we first initialize the weights of the backbone with those of the starting model (UDI), then we do continual SSL pre-training with an adapter (UniAdapter) on the target dataset by training only the adapter parameters. The obtained model is then used for finetuning for segmentation. We observe that GLARE reports consistent improvement for continual pre-training from UDI weights over all the datasets. On ADE20k, we obtain an improvement of +0.4+0.4 over UDI and on LoveDA, which is an out-of-domain dataset with respect to the original pre-training dataset of UDI (ImageNet (Deng et al., 2009)), we achieve an improvement of +0.6+0.6. This shows that GLARE is able to take advantage of existing encoded features and new data distribution to improve semantic understanding, which transfer in semantic segmentation.
Table 1: Comparison of SSL pre-trained models and continual pre-trained models starting from UDI (Su and Ji, 2024) on four semantic segmentation benchmarks. We report mIoU on the validation sets. We use the FPN (Kirillov et al., 2019) framework with 20k iterations and 2k for LoveDA. GLARE continual pre-training from UDI consistently shows improvements over the other pre-training strategies.
Method Backbone ADE20k P. Context Cityscapes LoveDA
@@ 6.2 Comparison with other continual pre-training methods
Table 2: Comparison of different continual pre-training framework. Experiments performed starting from UDI (Su and Ji, 2024) pre-trained on ImageNet (Deng et al., 2009). We show the segmentation performance of continual pre-trained models on LoveDA with a finetuning for 2k iterations.
@@ 6.3 Influence of the SSL backbone
An important question is the role of the SSL backbone used for the continual pre-training of GLARE. Specifically, we want to examine whether the performance gains observed when starting from UDI also hold when initializing from other SSL backbones, in particular FLSL (Su et al., 2023) and DINO (Caron et al., 2021), all pre-trained on ImageNet-1k (Deng et al., 2009). To investigate this, Table˜3 reports the results of continual pre-training with GLARE starting from these different backbones, using both HPT (Reed et al., 2022) and UniAdapter (Lu et al., 2023) frameworks. We use LoveDA (Wang et al., 2021a) as the target domain and evaluate performance after finetuning for semantic segmentation. We observe that GLARE consistently improves upon the original models. Moreover, starting from a stronger backbone (e.g., UDI) leads to larger gains. We further hypothesize that the original pre-training strategy influences how effectively GLARE can leverage a given backbone.
@@ 6.4 Do the continual pre-trained models forget?
Pre-training Pre-training Data ADE20k P. Context Cityscapes
UDI (Su and Ji, 2024) ImageNet (Deng et al., 2009) 41.1 49.2 74.7
@@ 6.5 Results in Classification
While our continual pre-training framework is primarily evaluated on semantic segmentation, it is not limited to this task. To demonstrate its broader applicability, we also apply GLARE continual pre-training to classification. Specifically, we experiment with two out-of-domain datasets: Derm7pt (Kawahara et al., 2019), a dermoscopic benchmark for skin lesion analysis containing $\sim$800 samples, and COVIDx (Wu et al., 2023), a large-scale x-ray benchmark for COVID-19 detection, where we subsample $20\%$ of the data for continual pre-training and finetuning to maintain a low-data setting. Models are initialized either randomly, from a DINO (Caron et al., 2021) pre-trained backbone, or from GLARE continual pre-training starting from DINO, and finally finetuning of a classification head. We report in Table˜5 the top-1 accuracy of models trained from scratch, initialized from DINO, or from DINO $\rightarrow$ GLARE. We observe that DINO $\rightarrow$ GLARE consistently outperforms the other initializations, demonstrating the effectiveness of GLARE continual pre-training even for tasks beyond semantic segmentation.
@@ 7.1 Different levels of understanding in GLARE
Table 6: Impact of the different pre-training levels in GLARE. We report the mIoU of finetuned continual pre-trained models when trained with different levels of GLARE on $20\%$ of ADE20k.
@@ Appendix A Details About Datasets
In this paper, we perform continual learning and report results on four semantic segmentation datasets: ADE20k (Zhou et al., 2017), Pascal Context (P. Context) (Mottaghi et al., 2014), Cityscapes (Citys.) (Cordts et al., 2016) and LoveDA (Wang et al., 2021a). We use the training set of these datasets to apply our continual learning setup with different pre-training methods such as UDI (Su and Ji, 2024), FLSL (Su et al., 2023), GLARE. We then finetune the resulting models for segmentation using FPN (Kirillov et al., 2019) and report the performances of the models on the validation sets. In the following we provide more details on these datasets.
@@
LoveDA (Wang et al., 2021a). This focuses on different geographical environments between urban and rural. It contains high spatial resolution (0.3m) remote sensing images containing objects at different scales, complex backgrounds and inconsistent class distributions. Its training set is composed of 2,522 images, its validation set is composed of 1,669 images and a test set of 1,796 images.
@@ C.1 Ablation of region sampling
Table 9: Ablation of the region sampling strategy. Experiments performed with UDI $\rightarrow$ GLARE with a finetuning of 2k iterations on LoveDA. $M$ represents the number of sampled regions considered.
@@ C.2 Effect of blurring strategy
random initialization 19.3 37.4 34.4
block-wise blurring 51.5 70.2 64.1
random blurring 51.5 70.2 64.3
@@ Appendix D Visualization of Attention Maps
In this section we visualize some attention maps from the last block of the ViT encoder, using the [CLS] token as the query token. Figure˜4 shows the attention from DINO (Caron et al., 2021), the original pre-trained weights from UDI, and a GLARE continual pre-trained model starting from UDI (Su and Ji, 2024) weights, finetuned on ADE20k. We observe similarities in how the attention is distributed across the images, focusing on various details such as foreground objects, object parts, and the background. GLARE continual pre-trained model demonstrates reduced noise relative to UDI, with its attention more precisely directed toward specific objects or regions. When using GLARE for continual pre-training, the model leverages what has been learned before and learns supplementary semantics specific to the dataset.